OpenStudy (anonymous):

determine whether the sequence converges or diverges: ln(n)/(n)^1/2... I have 2(n)^1/2 divided by n. But now what??

OpenStudy (experimentx):

simplify and take limit.

OpenStudy (anonymous):

Use L'Hospital's rule for \[ g(x) = \frac{\ln (x)} {\sqrt x} \]

OpenStudy (anonymous):

you get \[ \frac { \frac 1 x} {-\frac 1 { 2 \sqrt x}}= - \frac 2 {\sqrt x} \] which goes to zero a x go to Infinity.

OpenStudy (anonymous):

Oh yeah, I was wondering why you have 2/(x)^1/2 instead of \[2\sqrt{x} \div x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you undersatnd it now?

OpenStudy (anonymous):

I was asking why should I have \[2\div \sqrt{x} \] and not \[2\sqrt{x}\div x\]. They are different see?

OpenStudy (anonymous):

They are the same \[ \frac {2 x^{1/2}} {x}= \frac {2}{x^{1-\frac 12 }} = \frac {2}{\sqrt x} \]

OpenStudy (anonymous):

ok, thank you!!